the table lists the average annual cost (in dollars) of tuition and fees at a certain 2-year college for selected years, where year 1 represents 1992, year 3 represents 1994, and so on. the equation y= 31.33x + 1045 can be used to approximate the data. Predict the average annual cost in 2007 to the nearest dollar.

Year - Cost (in dollars)

1. - 1053

3. - 1139

5 -    1237

7 -    1287

9 - 1327
Answer #1

y = 31.33x + 1045

plug x = 2007 - 1991 = 16

y = 31.33 (16) + 1045

= 1546

Average annual cost in 2007 is $ 1546